OpenStudy (amorfide):

I know that e^(ix)=cosx+isinx but why does it? I'm bored of memorising things, can someone explain it so I can understand why?

OpenStudy (osprey):

is this de Moivre's theorem in complex numbers ? You ask "why" which is often a tricky q to ask in maths/science. After all i is the square root of minus one which only "exists" in the mind of the imaginary part of a a complex number. Actually, thinking about it, is this an equation or and identity ?????

OpenStudy (amorfide):

Identity I think demoivre's yes! although i think demoivres is to powers of n I just want to know why this specific example or in general if you can explain, is the same as this I want to start understanding math instead becoming a person who only memorises

OpenStudy (steve816):

Yeahh, pretty stumped with this as well, so is this what the equation is saying?\[\large e^{\pi \sqrt{-1}} = -1\]

OpenStudy (amorfide):

that is eulers identity I guess I am talking about demoivres formula

OpenStudy (steve816):

ah okee

zepdrix (zepdrix):

Define a complex value using polar coordinates,\[\large\rm z=\cos x+i \sin x\]Differentiate,\[\large\rm z'=-\sin x+i \cos x\]This step is a little tricky, factor an i out of each term,\[\large\rm z'=i(i \sin x+\cos x)=i(\cos x+i \sin x)\]Oh but cosx+isinx is what we started with! So we've determined this,\[\large\rm z'=i z\]Let's integrate and see what happens,\[\large\rm \frac{dz}{dx}=iz\qquad\to\qquad \int\limits\frac{dz}{z}=\int\limits i~dx\]Giving us,\[\large\rm \ln z=ix\]Exponentiating,\[\large\rm z=e^{ix}\]So we've shown that our complex number has two equivalent representations, cosx+isinx but also e^{ix}.

zepdrix (zepdrix):

That's one silly way to establish the relationship. You can also look at the uhhh Taylor Series for sine and cosine, and together they will relate to e^x Series in an interesting way :D

OpenStudy (amorfide):

I think starting at Z=cosx+isinx, which is the result we want to get to, is not a good way to start to explain something hmm the taylor series is definitely an interesting approach! Never even thought of that, could use e^x and let x=ix I suppose! thank you Zepdrix :) how do we know that integration works as naturally with imaginary numbers, as it does other numbers?

OpenStudy (eliesaab):

Also take the series of e^{ix} \[ 1+i x-\frac{x^2}{2}-\frac{i x^3}{6}+\frac{x^4}{24}+\frac{i x^5}{120}-\frac{x^6}{720}-\frac{i x^7}{5040}+\frac{x^8}{40320}+\frac{i x^9}{362880}-\frac{x^{10}}{3628800}+O\left(x^{11}\right) \] The even powers will give you cos(x) and i(odd powers)= i sin(x)

zepdrix (zepdrix):

Ahh that's a good question :D I'd have to look at my Complex Analysis book to remember how they justify integration techniques. I can't seem to remember.

OpenStudy (amorfide):

Ooo, any chance you can point me to the book online?

OpenStudy (eliesaab):

What do you mean by complex integration? Do you mean contour integrals?

OpenStudy (amorfide):

yes that is what I mean